Parth (parthkohli):

An interesting conjecture. Any even whole number more than or equal to 4 can be expressed as a sum of two prime numbers. For example: \[6=3+3\] \[32=19+13\] The list goes long but I am stopping right here.

OpenStudy (anonymous):

how about 2000 :P

Parth (parthkohli):

I don't really know it very long. It can be 991 + 1009, I think.

Parth (parthkohli):

Yes, it is.

OpenStudy (anonymous):

it is the property

OpenStudy (anonymous):

what about 2?

Parth (parthkohli):

1 + 1(because 1 is considered prime in some cases)

OpenStudy (anonymous):

1 is not prime no

OpenStudy (anonymous):

hmm...it is a matter of some debate

Parth (parthkohli):

I know, but it does have the properties of a prime number.

OpenStudy (anonymous):

it is not true 4 2

Parth (parthkohli):

Let's leave 2 behind. More numbers?

OpenStudy (anonymous):

@ParthKohli if you can solve this 5569652 it only proves that you have a calculator out there helping you :P lol

Parth (parthkohli):

I guessed the 2000 one lol

OpenStudy (anonymous):

since it is satiesfying aa even no it is property

OpenStudy (unklerhaukus):

one is an even whole number that cannot be expressed as a sum to two prime numbers

Parth (parthkohli):

One is odd* @UnkleRhaukus

OpenStudy (unklerhaukus):

odd, hmm yeah i didn't really think about what this word meant haha

OpenStudy (anonymous):

uncle pagla gaya h

OpenStudy (anonymous):

isnt this Goldbach's Conjecture? its for even numbers greater than or equal to 4, so dont worry about 2.

OpenStudy (anonymous):

parth qn change karo

OpenStudy (anonymous):

For conventional purposes, 1 is not a prime, because we want integers to have a unique prime factorization. If 1 was considered a prime, then we would have:\[20=2^2\cdot 5=1\cdot 2^2\cdot 5=1^{100}\cdot 2^2\cdot 5=\cdots\]

OpenStudy (unklerhaukus):

is zero odd even or neither?

OpenStudy (anonymous):

Even.

OpenStudy (anonymous):

bhai joe see property of prime no see defination of prime

OpenStudy (anonymous):

Why is 0 even?

OpenStudy (anonymous):

because it is a multiple of 2

OpenStudy (anonymous):

2*0 = 0;

OpenStudy (anonymous):

3*0=0

Parth (parthkohli):

Yes, this is Goldbach's

OpenStudy (anonymous):

because it can be written as:\[0=2(0)\]any even number can be written as:\[n=2k\]any odd number can be written as:\[n=2k+1\]

OpenStudy (anonymous):

O is even because it is divisible by 2.

OpenStudy (anonymous):

oh i see

Parth (parthkohli):

I changed it.

OpenStudy (anonymous):

even comes after odd and before odd no -1 0 1

OpenStudy (unklerhaukus):

ok i'm convinced zero is even Can we use the Goldbach's Conjecture to find prime numbers?

Parth (parthkohli):

We cannot.